The record of births, over the last 100 years maintained by the municipal council of a town showed that 200 children were born blind during that period. On the assumption that the number of children born blind in an year follows Poisson distribution, estimate the number of years in which there were (a) no blind births (b) one blind birth (c) two blind births 


1
Expert's answer
2021-11-09T16:13:04-0500

Let "X=" the number of births in year: "X\sim Po(\lambda)"



"\lambda=200/100=2"

(a)

"P(X=0)=\dfrac{e^{-2}(2)^0}{0!}=e^{-2}\approx0.135335"

"0.135335(100)=14(years)"

(b)

"P(X=1)=\dfrac{e^{-2}(2)^1}{1!}=2e^{-2}\approx0.270671"

"0.270671(100)=27(years)"

(c)

"P(X=2)=\dfrac{e^{-2}(2)^2}{2!}=2e^{-2}\approx0.270671"

"0.270671(100)=27(years)"